OpenStudy (nerdsarecool):

Jose had 2/7 acres of land. He planted 4/5 of it with corn. How many acres did he plant with corn?

OpenStudy (texaschic101):

4/5 of 2/7 = 4/5 * 2/7 = 8/35 acres